the perimeter of a rectangle is 46 cm. the width is 5 more than twice the length. what is the area of the rectangle
Answer by stanbon(75887)

the perimeter of a rectangle is 46 cm. the width is 5 more than twice the length. what is the area of the rectangle
----------------
Let the length be "x".
Then the width = "2x+5"
-----------------------------
Equation:
Perimeter = 2(W + L)
46 = 2(2x+5 + x)
23 = 3x+5
3x = 18
x = 6 cm (length)
2x+5 = 2*6+5 = 17 cm (width)
===============================
Cheers,
Stan H.